### v4

# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

Abstract
Classical federated learning (FL) assumes that the clients have a limited amount of noisy data with which they voluntarily participate and contribute towards learning a global, more accurate model in a principled manner. The learning happens in a distributed fashion without sharing the data with the center. However, these methods do not consider the incentive of an agent for participating and contributing to the process, given that data collection and running a distributed algorithm is costly for the clients. The question of rationality of contribution has been asked recently in the literature and some results exist that consider this problem. This paper addresses the question of simultaneous parameter learning and incentivizing contribution in a truthful manner, which distinguishes it from the extant literature. Our first mechanism incentivizes each client to contribute to the FL process at a Nash equilibrium and simultaneously learn the model parameters. We also ensure that agents are incentivized to truthfully reveal information in the intermediate stages of the algorithm. However, this equilibrium outcome can be away from the optimal, where clients contribute with their full data and the algorithm learns the optimal parameters. We propose a second mechanism that, in addition, enables the full data contribution with certain tradeoffs between budget balance and truthfulness. Large-scale experiments with real (federated) datasets (CIFAR-10, FEMNIST, and Twitter) show that these algorithms converge quite fast in practice, yield good welfare guarantees and better model performance for all agents.
Papers/Algorithms NE Learning Guarantee Payments Reported Quantity Truthful Data Contribution Data Privacy
Procaccia et al. (2025) Existence ✓($B$) Distrib. DSIC Partial
Bornstein et al. (2024) Existence Cost DSIC Full
Blum et al. (2021) Existence Contrib. Partial
Murhekar et al. (2023) Convergence ✓($B$) Contrib. + Cost Partial
Karimireddy et al. (2022) Existence Contrib. + Cost Full
This paper (UPBReD) Convergence AC Contrib.+ Gradient Partial
This paper (T-UPBReD) Convergence AC Contrib.+ Gradient EPIC Partial
This paper (2P-UPBReD) Convergence AC ($B$) Contrib.+ Gradient Full
Table 1: Comparative analysis of 2P-UPBReD against existing incentive mechanisms in FL. NE indicates whether the existence or best response convergence to a Nash Equilibrium is established. Learning Guarantees specifies the nature of the model convergence (e.g. Algorithmic Convergence (AC)). Payments denotes the use of monetary transfers and whether they are Budget Balanced ($B$). Reported Quantity identifies the strategic variables (e.g., distribution, contribution, cost, or gradients). Truthful indicates the formal incentive properties (e.g., DSIC, EPIC) with respect to the reported quantity. Data Contribution specifies the amount of data contributed by an agent at equilibrium (Partial vs. Total). Data Privacy specifies if datasets are private and not shared with the center. Unlike prior works, our mechanisms provide joint convergence guarantees for both model weights and strategic gradients.
1 Introduction
A high-quality machine learning model is built when the model is trained on a large amount of data. However, in various practical situations, e.g., for languages, images, disease modeling, such data is split across multiple entities. Moreover, in many applications, data lies in edge devices, and a model is learnt when these edge devices interact with a parameter server. Federated Learning (FL) Konečnỳ (2016); McMahan et al. (2017); Kairouz et al. (2021); Zhang et al. (2021) is a recently developed distributed learning paradigm where the edge devices are users’ personal devices (like mobile phones and laptops) and FL aims to leverage on-device intelligence. In FL, there is a center (also known as the parameter server) and several agents (edge devices). To learn the model parameters, the edge devices can only communicate through the center. However, a major challenge of FL is data-scarcity, i.e., owing to its limited storage capacity, each edge device possesses only a limited amount of data, which may not be sufficient for the learning task. Hence, the end user participates in a distributed learning process, where they exploit the data of similar users present in the system. As a result of their participation, the users gain some benefit from the shared model that is learned and incur some loss for their contributions to the learning process.
In the classical federated learning problem, it is typically assumed that the agents participate with all of their data points voluntarily. However, in the presence of rational users, such an assumption may not always hold since sampling data is costly for agents. It may happen that some agents try to contribute very few data-points and try to exploit the system by learning based on others’ data-points. This phenomenon is called free-riding Karimireddy et al. (2022) and disincentivizes honest participants to contribute their data to the FL process. Hence, naturally, a few recent works have concentrated on designing incentive mechanisms in FL, Murhekar et al. (2023); Donahue and Kleinberg (2021); Blum et al. (2021).
One of the most successful use-cases of FL is in hospital management systems. For instance, in a given geographical region every hospital can sample medical records from its patients for disease modeling. To learn the model reliably, each hospital needs to collect a large amount of data, which is generally very expensive. Instead, they might participate in a federated learning (FL) process where different hospitals can train the model locally on their individual datasets but update a consolidated global model parameter. Given that different hospitals have different capabilities in collecting data, it may be best for certain hospitals not to sample any data and therefore not incur any cost, yet get the learning parameters from the FL process if there is a sufficient population in the FL who are sharing their learned parameters. This occurrence of free-riding is not ideal, and an important question to ask is:
“Can a mechanism be designed that incentivises each user to participate and contribute their maximum data in FL and simultaneously learn the optimal model parameters?”
In this paper, we address this question in two stages. First, we consider a utility model of the agents where the quality of the learned parameter and the contribution levels of all the agents give every agent some benefit, while their individual contribution levels lead to their personal costs. We propose an algorithm, namely Updated Parameter Best Response Dynamics (UPBReD) in the FL setting that achieves simultaneous contribution and learning of the model parameters by the agents. We introduce payments to ensure that agents truthfully share their contributions. However, this method suffers from a sub-optimal contribution by the agents and therefore the global learning is also sub-optimal. In the second stage, we improve this dynamics with monetary transfers such that it incentivizes all agents to contribute their maximum amount of data to the learning process. This mechanism, namely Two Phase Updated Parameter Best Response Dynamics (2P-UPBReD), learns the optimal parameters for all agents and asks the data-consumers to pay and pays the data-contributors, without keeping any surplus. 2P-UPBReD is distinguished in the fact that it simultaneously learns the optimal model parameters, incentivizes full contribution from the agents, yet is quite simple to use in practice.
1.1 Our contributions
Our contributions can be summarized as follows.
  • We propose UPBReD (Algorithm˜1) that allows simultaneous learning and contribution by the agents in FL, and show that it converges to a pure strategy Nash equilibrium (Theorem˜1).
  • We add payments to UPBReD to construct a mechanism T-UPBReD that elicits truthful reports from agents in every round (Theorem˜2).
  • Since the Nash equilibrium can be different from the socially optimal outcome (where all agents contribute their $s^{\max}$) (˜1), we propose an updated and cleaner two-phase mechanism, 2P-UPBReD (Algorithm˜2) that allows monetary transfers, and is budget balanced. The data contributors (contributing above average quantities of data) get paid, and the data consumers (contributing below average quantities of data) make the payments. However, both types of agents learn the optimal model parameters.
  • We show that the Nash equilibrium for 2P-UPBReD now shifts to where all agents make their maximum contribution of data $s^{\max}$ and learn the optimal model parameters $w^{\texttt{{OPT}}}$ (Theorem˜3).
  • Experiments on real datasets (CIFAR-10, FEMNIST, and Twitter) demonstrate that while UPBReD leads to suboptimal social welfare, 2P-UPBReD achieves performance comparable to FedAvg for all datasets (see Section˜5). We introduce FedAvgStrategic to characterize the incentive gap in FL. Our results demonstrate standard aggregation methods (FedAvg) without incentives are highly susceptible to strategic agents who choose to withhold data, leading to a significant degradation in global model quality (Figure˜3). We also compare the algorithms on non-iid datasets partitioned using the Flower framework (Beutel et al., 2020).
We compare the results of this work with other closely related work in incentives for FL in Table˜1.
1.2 Related Work
Federated learning Konečnỳ (2016) has gained significant attention in the last decade or so. The success story of FL is primarily attributed to the celebrated FedAvg algorithm McMahan et al. (2017), where one of the major challenges of FL, namely communication cost is reduced by local steps. In subsequent works, several other challenges of FL, such as data heterogeneity Karimireddy et al. (2020b); Ghosh et al. (2020a), byzantine robustness Yin et al. (2018); Karimireddy et al. (2020a); Ghosh et al. (2021), communication overhead Stich et al. (2018); Karimireddy et al. (2019); Ghosh et al. (2020b) and privacy Wei et al. (2020); Truex et al. (2020); Kumar et al. (2025) were addressed.
Later works focused on incentivizing client participation with advanced aggregation methods. Approaches included dynamic weighting of client updates to ensure that the global model outperforms local models Cho et al. (2022), Shapley value to assess contributions Tastan et al. (2024), contract theory to maximize fairness in data contribution Karimireddy et al. (2022), and fairness by eliminating malicious clients and rewarding those who enhance performance of the model Gao et al. (2021),  Procaccia et al. (2025) aim to maximize utilitarian welfare in the heterogenous setting under PAC learning constraints.
Monetary incentives in FL have been explored through various mechanisms. Yu et al. (2020) proposed dynamic budget allocation to maximize utility and reduce inequality. Blockchain-based methods Pandey et al. (2022) incentivize high-quality data contributions under budget constraints, and client rewards have been shown to improve final model utility Yang et al. (2023)Liu et al. (2020) use a blockchain based system to enable Shapley value based profit distributions. Georgoulaki and Kollias (2023) analyzed utility-sharing in FL games, showing a price of anarchy of two and price of stability of one under budget-balanced payments. Auction-based FL systems use reinforcement learning to let agents adjust bids for profit while preserving accuracy Tang and Yu (2023), and centers can optimize budgets to enhance utility and reduce delays Tang and Yu (2024).
Coalitional approaches have also been studied. Donahue and Kleinberg (2021) modeled FL as a coalitional game allowing joint training, while Ray Chaudhury et al. (2022) introduced CoreFed to ensure core stability. Their work was generalized to ordinal utility settings Chaudhury et al. (2024), and Shapley value-based approaches ensure reciprocal fairness Murhekar et al. (2024). However, these do not consider non-cooperative agents choosing data contributions strategically. Repeated-game formulations Mao et al. (2024) show that subgame perfect equilibria can be inefficient, prompting budget-balanced mechanisms that ensure social efficiency and individual rationality. Truthful cost reports are explored by Bornstein et al. (2024). Chakarov et al. (2025) introduce a budget balanced payment rule that is Bayesian incentive compatible in the heterogenous setting. Incentive-aware learning frameworks by Blum et al. (2021) and their extension with best response dynamics and budget-balance by Murhekar et al. (2023) achieve Nash equilibria and maximum welfare under constraints. Yet, these works overlook simultaneous model training, a gap our work addresses by incorporating strategic data contribution and model parameter learning together.
2 Preliminaries
Consider an FL setup where a set of data contributors, given by $N=\{1,2,\ldots,n\}$, is interacting with a center. Each data contributor (agent) $i$ has access to a private labeled dataset $D_{i}$, with the size of the dataset given by $s_{i}^{\max}=|D_{i}|$. The agents are interested in learning a parameter vector $w\in\mathbb{R}^{m}$ from these data so that it helps them predict some unlabeled data accurately (e.g., to perform a classification task). However, each individual agent has a limited amount of data, and learning $w$ only from that data may not be accurate enough. So, they learn this parameter via a federated learner such that the model is trained on the consolidated data of all the $n$ users. Assume that the datasets are drawn from the same distribution, e.g., all the agents sample human disease data from a certain geographical location. However, sampling such data is costly, and agent $i$ incurs a cost, given by a function $c_{i}(s_{i})$, when it trains the model locally on its chosen dataset of size $s_{i}\in S_{i}:=[0,s_{i}^{\max}]$. Along with these costs, there may be some other arbitrary costs that an agent incurs based on the mechanism and other agent contributions. Under this setup, each agent $i$ gets a utility based on how much data $s$ has been chosen by the agents to sample and train on. Hence, the utility is given by
$u_{i}(w,s_{i},s_{-i})=v_{i}(w,s_{i},s_{-i})-z_{i}(s_{i},s_{-i}),$ (1)
where $s_{-i}$ is the data chosen by the agents other than $i$ in this FL process. The function $v_{i}(w,s_{i},s_{-i})$ is the valuation function, which denotes the benefit to agent $i$ if the parameter learned by the center is $w$ and the agents contribute by running the FL algorithm on their dataset sizes given by the vector $s=(s_{i},s_{-i})$. The function $z_{i}(s_{i},s_{-i})$ is the effective cost to agent $i$ which may depend on the data contributions of all agents.
FL is an iterative process, the center updates a parameter $w^{t}$ at round $t$, agents use this to choose their dataset size $s_{i}^{t+1}$, compute and share gradients $d_{i}^{t+1}\in\mathbb{R}^{m}$ with the center. The center aggregates gradients to update the parameter to $w^{t+1}$. We use the superscript $t$ when referring to these terms in a specific iteration of the training round.
Remark (Arguments of the $u_{i}$ function). Note that we have explicitly assumed that the utility function $u_{i}$ depends on the parameter learned and the data contributions $(w,s_{i},s_{-i})$. We now motivate this dependence. In FL problems, we typically run iterative algorithms to optimize the accuracy (or loss) function. Note that, with this, the weight $w^{t}$ at time $t$ depends on all the $s_{i}^{\ell}$, $\ell<t,\forall i\in N$. However, the utility $u_{i}$ at time $t$ depends on $(s_{i}^{t},s_{-i}^{t})$ as well, which is not captured through $w^{t}$. Hence, we require the said explicit dependencies. We represent this dependency via two functions: the valuation function $v_{i}$ gives the benefit the agent gets from the parameter learned and the data contributions, while the effective cost is only dependent on all players’ data contributions.
Remark (Realizing the $v_{i}$ function). $v_{i}(w,s)$ denotes the benefit agent $i$ derives from the learned parameters at a given contribution vector. It can be realistically realised by an agent using the negative of loss computed on their data for the given parameter $w$. For experiments in Section˜5, we use the cross-entropy loss.
Since the agents are strategic, every agent $i$’s aim is to maximize its utility by appropriately choosing its strategy $s_{i}$ given the strategies $s_{-i}$ of the other players and the parameter $w$ chosen by the center. The center, on the other hand, is interested in learning the optimal parameter $w$ that maximizes the sum of the valuations as follows, when all agents contribute their maximum data-sizes, i.e., $s_{i}^{\max},i\in N$.
$w^{\texttt{{OPT}}}\in\operatorname*{\arg\!\max}_{w}\textstyle\sum_{i\in N}v_{i}(w,s^{\max})$ (2)
Note that the goal of the center does not consider the effective costs of the agents since those are incurred by the agents. We will refer to the term $\sum_{i\in N}v_{i}(w,s_{i},s_{-i})$ as the social welfare in this context. We are interested in the question of whether we can design an FL algorithm that can make $(s_{i}^{\max},s_{-i}^{\max})$ a Nash equilibrium of the underlying game, and the center can learn $w^{\texttt{{OPT}}}$. In this context, the Nash equilibrium is defined as follows.
Definition 1 (Nash equilibrium).
A pure strategy Nash equilibrium (PSNE) for a given parameter $w$ is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ of the agents such that $u_{i}(w,s_{i}^{*},s_{-i}^{*})\geqslant u_{i}(w,s_{i},s_{-i}^{*}),\ \forall s_{i}\in S_{i},\forall i\in N$.
This definition is a modification of the standard definition of Nash equilibrium since the equilibrium profile $(s_{i}^{*},s_{-i}^{*})$ depends on the parameter $w$ (we do not write it explicitly for notational cleanliness). We will refer to the term PSNE as Nash equilibrium (NE) in the rest of the paper.
We also aim for the following property of a mechanism that involves monetary transfers.
Definition 2 (Budget balance).
A mechanism that uses monetary transfers $p_{i}(s_{i},s_{-i})$ for every $s_{i}\in S_{i},i\in N$ is called budget balanced (BB) if $\sum_{i\in N}p_{i}(s_{i},s_{-i})=0$.
This property ensures that the net monetary in or out-flow is zero and the mechanism only allows monetary redistribution among the agents.
In our algorithms the agents are asked to share $s_{i}^{t}$ and $d_{i}^{t}$ in each round, this tuple $\theta_{i}^{t}=(s_{i}^{t},d_{i}^{t})$ forms the true type of an agent. The reported type profile $\theta=(\theta_{1},\ldots,\theta_{n})\in\Theta$ is then used to determine the global parameter $w^{t}=a(\theta^{t})$, where $a:\Theta\rightarrow\mathbb{R}^{m}$ is the decision rule implemented by the center. While agents strategically choose $s_{i}^{t}$ and the corresponding $d_{i}^{t}$ to maximise their utility they need not report this type truthfully. Utilities under misreported type profiles use both the reported and true types, when $\theta,v$ are the true type and valuation profiles and $\hat{\theta},\hat{v}$ are the reported type and valuation profiles the utility is computed as $u_{i}(\hat{\theta},\hat{v}|\theta,v)=v_{i}(a(\hat{\theta}),s)+z_{i}(s_{i},\hat{s}_{-i})$. Our truthfulness guarantee in this setting is ex-post incentive compatibility (EPIC), defined as follows.
Definition 3 (Ex-Post Incentive Compatible (EPIC)).
A mechanism with a given allocation rule $a$ and a payment rule $p$ is EPIC at round $t$ if for every agent $i\in N$ for the true type and valuation profile $\theta,v$ and a misreported type and valuation profile $\hat{\theta},\hat{v}$,
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})\leqslant u_{i}({\theta}^{t},{v}^{t}|\theta^{t},v^{t}).$
EPIC ensures that an agent gains no improvement in utility for misreporting when other agents are reporting truthfully. This is the strongest truthfulness notion for mechanism design with interdependent valuations (valuation of an agent depends on the types of all agents) since satisfying DSIC is impossible in such settings (see Mezzetti (2004) for further details). In the following section, we consider the general case where the effective cost is arbitrary.
3 Learning with arbitrary effective costs
In this section, we consider the effective cost of agent $i$, i.e., $z_{i}(s_{i},s_{-i})$, to be an arbitrary function of $s=(s_{i},s_{-i})$. We know that an alternative interpretation of NE (Definition˜1) is a strategy profile $(s_{i}^{*},s_{-i}^{*})$ where every agent’s best response to the strategies of the other players is its own strategy in that profile (see  (Maschler et al., 2020, e.g.)), i.e., $s_{i}^{*}\in\operatorname*{\arg\!\max}_{s_{i}\in S_{i}}\ u_{i}(w,s_{i},s_{-i}^{*}),\ \forall i\in N$. Hence, an algorithm that simultaneously updates all agents’ strategies with the best responses to the current strategies of the other players is called a best response dynamics of a strategic form game (see  (Fudenberg, 1991, e.g.) for a detailed description). In our problem, this approach cannot be directly employed since the center also needs to learn and update the model parameters $w$ as the agents choose their data contributions. We, therefore, propose a mechanism for FL that simultaneously updates both the agents’ strategies and the center’s choice of $w$. FL protocols iteratively use local gradients to update a global model, the mechanism we describe below captures strategic behaviour at every iteration by allowing agents to update their strategy (data set sizes used for gradient computation) at every realisation of the global model based on Equation˜1.
In this mechanism, each agent $i$ starts with some initial choices of $s_{i}^{0}$ and shares that with the center. [^1] Note that, only the number of data points, $s_{i}$ is shared with the center and not the data, which is consistent with the principle of FL. The center initializes a $w^{0}$ and broadcasts that and the entire initial data-contribution vector $s^{0}$ to all the agents. In every subsequent iteration $t$, each agent $i$ locally computes two quantities: (i) its updated contribution $s_{i}^{t+1}$ by taking one gradient ascent step w.r.t. its own contribution $s_{i}$, and (ii) the local gradient of agent $i$’s valuation component of the social welfare w.r.t. $w$. Both are evaluated at the current values of $w^{t}$ and $s^{t}$ and sent back to the center. The center calculates an updated $w^{t+1}$ that averages (in spirit of the FedAvg algorithm (McMahan et al., 2017)) all the local gradients sent by the agents. The center then shares $w^{t+1}$ and $s^{t+1}$ with all the agents. Formally, the updates are given as follows.
$s^{t+1}=s^{t}+\gamma g(w^{t},s^{t},\mu^{t}),$ $\qquad w^{t+1}=w^{t}+\eta\tilde{g}(w^{t},s^{t}),$ (3)
[…]
The mechanism is detailed out in Algorithm˜1. Our main result of this section is that under certain bounded derivative conditions, Algorithm˜1 always converges to a Nash equilibrium. We need a few matrices, defined as follows for every $w$ and $s$. For $i,j\in N$ and $k,\ell\in\{1,\ldots,m\},G(w,s)_{ij}=\frac{\partial^{2}}{\partial s_{j}\partial s_{i}}u_{i}(w,s);\tilde{G}(w,s)_{k\ell}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial w_{\ell}\partial w_{k}}v_{i}(w,s);H(w,s)_{ik}=\frac{\partial^{2}}{\partial w_{k}\partial s_{i}}u_{i}(w,s)$; $\tilde{H}(w,s)_{kj}=\frac{1}{n}\sum_{i\in N}\frac{\partial^{2}}{\partial s_{j}\partial w_{k}}v_{i}(w,s)$.
Assumption 1.
Consider the utility functions given by Equation˜1 where functions $v_{i},z_{i},i\in N$ are such that the following properties hold for every $w\in\mathbb{R}^{m}$ and $s\in\prod_{i\in N}S_{i}$, $\forall i,j\in N,\forall k,\ell\in\{1,\ldots,m\}$ (the matrices below are as defined above).
  1. 1.
    The matrices $G(w,s)+\lambda\mathbb{I}$ and $\tilde{G}(w,s)+\tilde{\lambda}\mathbb{I}$ are negative semi-definite.
  2. 2.
    We assume the following bounds: (i) $|G(w,s)_{ij}|\leqslant L$, (ii) $|\tilde{G}(w,s)_{k\ell}|\leqslant\tilde{L}$, (iii) $\|H(w,s)\|_{op}\leqslant P$, where $\|A\|_{op}:=\inf\{c\geqslant 0:\|Av\|\leqslant c\|v\|,\forall v\}$, (iv) $\|\tilde{H}(w,s)\|_{op}\leqslant\tilde{P}$.
Note that the definition of $G$ and $\tilde{G}$ do not imply the strong concavity of $u_{i}$ or $v_{i}$. Moreover, the above assumptions are standard and have appeared in the literature before (see Murhekar et al. (2023)). We define the following expressions for a cleaner presentation. $W_{1}=\sqrt{1+\gamma^{2}n^{2}L^{2}-2\gamma\lambda}\sqrt{\tilde{P}^{2}\gamma^{2}},W_{2}=\sqrt{1+\eta^{2}m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda}}+\sqrt{P^{2}\eta^{2}},E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2},T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{E}{\epsilon}}\right)}}{{\left(\ln{\frac{1}{W}}\right)}},$ where $w^{0}\in\mathbb{R}^{m},s^{0}\in\prod_{i\in N}S_{i}$ and $W=\max\{W_{1},W_{2}\}$. We are now ready to present the main result of this section. Full proofs of the technical results are presented in the Appendix.
Theorem 1 (UPBReD convergence to NE).
Under ˜1, UPBReD (Algorithm˜1) converges to a Nash equilibrium. Formally, for a given $\epsilon>0$, for every initial value $(w^{0},s^{0})$ of Algorithm˜1, the gradients $\|g(w^{T},s^{T},\mu^{T})\|<\epsilon$ and $\|\tilde{g}(w^{T},s^{T})\|<\epsilon$, for all $T\geqslant T_{0}(w^{0},s^{0})$, when the step sizes are chosen as follows: $\gamma<\min\left\{1,\frac{1}{\tilde{P}},\frac{2\lambda}{n^{2}L^{2}},\frac{\lambda-\tilde{P}}{n^{2}L^{2}-\tilde{P}^{2}}\right\},\text{ given }\lambda>\tilde{P}$, and $\eta<\min\left\{1,\frac{1}{P},\frac{2\tilde{\lambda}}{m^{2}\tilde{L}^{2}},\frac{\tilde{\lambda}-P}{m^{2}\tilde{L}^{2}-P^{2}}\right\},\text{ given }\tilde{\lambda}>P$.
1
Input: Step size $\gamma,\eta$, initialization $w^{0}$, $s_{i}^{0}$ for $i\in N$, number of iterations $T$
Output: $w^{T}$
2
3for $t=0$ to T1T-1 do
4 Center broadcasts $w^{t},s^{t}$
5
6 foreach agent $i\in N$ in parallel do
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
8      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
9      Send $s_{i}^{t+1},d_{i}^{t+1}$ to the center
10
11
12 Center updates: $w^{t+1}=w^{t}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t+1}=w^{t}+\eta\tilde{g}(w^{t},s^{t})$
13
14
15return $w^{T}$
16
Algorithm 1 Updated Parameter Best Response Dynamics (UPBReD)
The theorem provides a guarantee of convergence of Algorithm˜1 for any arbitrary initial condition $(w^{0},s^{0})$. It needs to run for a minimum number of iterations $T_{0}(w^{0},s^{0})$, with appropriate parameters of $\gamma$ and $\eta$ that govern the gradient ascent rates of the agents’ and center’s objectives respectively. Note that for a better convergence, i.e., a smaller $\epsilon$, the algorithm needs to run longer. The parameter $W$ determines the contraction factor of the recurrence of $\|g(w^{t},s^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}$ and is chosen to be smaller than unity by choosing $\gamma$ and $\eta$ appropriately.
Figure˜1 shows feasible regions of $\gamma$ for certain $n,L,\lambda,\tilde{P}$.
(a) $\tilde{P}=1$
(b) $\lambda=1.1$
Figure 1: Shaded regions show the feasible choices of $\gamma$ for $n=2,L=1$ and $\tilde{P}$ and $\lambda$ as shown. The dashed lines show the boundary of the regions in the legends. A similar set of choices is true for $\eta$.
Remark. The step sizes $\gamma$ and $\nu$ in our analysis scale inversely with $m$ and $n$, and certain restrictions on $\lambda$ and $\tilde{\lambda}$ may appear restrictive. In Section 4, we address these issues and propose a learning algorithm with step sizes independent of $m$ and $n$.
In practice, however, we observe that sufficiently small constant values of $\gamma$ and $\nu$ work well, indicating that the scaling requirement in the theorem is merely a theoretical artifact rather than a practical limitation.
3.1 Application of Theorem˜1: effective costs
are only personal
One special but practical case of the above setup is when the effective cost is only borne by the agent. Mathematically, it is represented as $z_{i}(s_{i},s_{-i})=c_{i}(s_{i}),\forall s_{i}\in S_{i},s_{-i}\in S_{-i},\forall i\in N$. We assume convex cost functions $c_{i},\forall i\in N$, which is a commonly used assumption Li and Raghunathan (2014) in the literature. This setup is consistent with the assumptions of ˜1, and according to Theorem˜1, Algorithm˜1 converges to a Nash equilibrium, which we denote as $(w^{*},s^{*})$. However, it is possible that none of the following things happen: (i) $s^{*}\neq s^{\max}$, i.e., the agents do not contribute their entire data for the federated learning process, leading to a suboptimal learning, or (ii) $w^{*}\neq w^{\texttt{{OPT}}}$ (see Equation˜2 for the definition), which is neither the objective of the center nor the agents. The following example shows such an instance.
Example 1 (NE different from socially optimal).
Consider the federated learning setup with two agents that have linear costs and are trying to learn a model with two parameters. The valuation function is identical and is given by $v_{i}(w,s)=1-\frac{(1-w_{1})^{2}+(2-w_{2})^{2}}{s_{1}+s_{2}},i=1,2$. The cost functions for agents $1$ and $2$ are given by $c_{1}(s_{1})=0.04\cdot s_{1}$ and $c_{2}(s_{2})=0.02\cdot s_{2}$ respectively. The agents’ strategy sets are $S_{i}=[0,5],i=1,2$. Notice that, at $s=s^{\max}=(5,5)$, the value of $w^{\texttt{{OPT}}}$ (see Equation˜2) is $(1,2)$. It yields an optimal social welfare of $2$. However, this is not an NE, since the derivatives $\partial u_{i}/\partial s_{i}|_{s_{i}^{\max}}<0$ for both $i=1,2$. Hence, the best response of each agent is to reduce $s_{i}$. However, from this point, with the choice of $\gamma=0.25$ and $\eta=0.25$, Algorithm˜1 converges to an NE profile of $w^{*}=(0.5,1.5),s^{*}=(0,5)$ that yields a social welfare of $1.8$. ∎
So, our objective in this paper is to allow monetary transfer among the agents so that we get the best of both worlds: (a) We achieve convergence to a Nash equilibrium of $(w^{\texttt{{OPT}}},s^{\max})$, where all agents to contribute their entire data and center learns the optimal parameter, and (b) the transfers are budget balanced, i.e., the center does not accumulate any money – it is used just to realign the agents’ utilities to reach the desired Nash equilibrium. We discuss this mechanism in Section˜4.
While agents strategically choose $s_{i}$ to maximize their utility, we assume truthful reporting of these choices. In UPBReD, agents report $s_{i}^{t}$ and $d_{i}^{t}$ in each round. Section˜A.2 shows an example where misreporting can improve utility. The next subsection introduces payments to ensure truthful reporting of $s_{i}^{t}$ and $d_{i}^{t}$.
3.2 Truthful elicitation with payment rules
So far, we have assumed that the agents report their type truthfully in every round, i.e. if an agent is training on $s_{i}^{t}$ amount of data at time $t$ the agent reports $\theta_{i}=(s_{i}^{t},d_{i}^{t})$ truthfully. However, an agent can misreport their strategy and corresponding gradient update to improve their utility as observed in Section˜A.2.
A well established goal in social choice theory is to choose an allocation that maximizes the social welfare (Green and Laffont, 1977). In our setting this maximizing allocation for a given type profile $\theta^{t}$ at round $t$ will be $a_{*}({\theta}^{t})=\operatorname*{\arg\!\max}_{w\in\mathbb{R}^{m}}{\sum_{i\in N}v_{i}(w,s^{t}})$
However, in our FL setting, we cannot compute this welfare maximizing allocation. Instead, we use $a_{w}^{t}$, which iteratively takes gradient ascent steps towards the welfare maximizing allocation as shown in Algorithm˜1. The initial allocation is chosen as $w^{0}$. And future allocations are computed iteratively as $a_{w}^{t}(\theta^{t})=w^{t}(s^{t},d^{t})=w^{t-1}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t}$. Due to this difference, we need an additional assumption for truthfulness. We assume the difference between the welfare at our computed allocation and that of the maximum at truthful reports is bounded by a constant.
Assumption 2.
$\sum_{i\in N}v_{i}(a_{*}(\theta^{t}),s^{t})-\sum_{i\in N}v_{i}(a_{w}^{t}(\theta^{t}),s^{t})<\Gamma,\>\forall\theta^{t}\in\Theta,\ \forall t=1,\ldots,T$.
Remark. Note that the above assumption is weaker than that of bounded valuations, that are standard when choosing allocation that are welfare maximizing. When valuations are non-negative and bounded above by some constant $M$, $\Gamma$ can be upper-bounded by $n.M$. Our mechanism remains stable for any $\Gamma^{\prime}\geq\Gamma=n.M$, ensuring that the center only needs a coarse estimate of the maximum valuation to satisfy ˜2.
Consider the following payment rule:
$p_{i}^{*}(\hat{\theta}^{t},\hat{v}^{t})=\sum_{j\neq i}\hat{v}_{j}^{t}-\phi(\hat{v}_{i}^{t},v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))-h_{i}(\hat{\theta}_{-i}^{t})$ (4)
where, $\phi=0\ \text{if }\hat{v}^{t}_{i}=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t});\phi=\Gamma+(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{s}^{t}))^{2}\>\text{ otherwise}$.
T-UPBReD initializes a parameter $w^{0}$ and agents begin with an initial vector $s^{0}$. In every subsequent round T-UPBReD follows the following steps at round $t$.
  1. 1.
    The center computes and shares a parameter $w^{t-1}$ computed in the previous round with the agents.
  2. 2.
    Agents compute the gradients $d_{i}^{t}$ with respect to $w^{t-1}$ on their chosen dataset size $s_{i}^{t-1}$ then strategically pick their dataset size for the next round $s_{i}^{t}$ using Equation˜3, which forms the true type of the agent $\theta^{t}_{i}=(s_{i}^{t},d_{i}^{t})$.
  3. 3.
    Agents share their type as $\hat{\theta}_{i}^{t}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$ which can be different from their true type. The center computes an allocation using the reported types using $a_{w}^{t}$ consistent with Equation˜3.
  4. 4.
[…]
  5. 5.
    Agents report their valuation as $\hat{v}_{i}^{t}$ which may be different from the true valuations. The center then computes payments $p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$ for each agent.
With the given allocation rule $a_{w}^{t}$ and payment rule $p$ described in Equation˜4 for a true type and valuation profile $\theta^{t},v^{t}$, a reported type and valuation profile $\hat{\theta}^{t},\hat{v}^{t}$, the utility realised by agent $i$ at round $t$ is given by $u_{i}(\hat{\theta}^{t},\hat{v}^{t}|\theta^{t},v^{t})=v_{i}(a_{w}^{t}(\hat{\theta}^{t}),s^{t})-z_{i}(s_{i}^{t},\hat{s}_{-i}^{t})+p_{i}^{*}(\hat{\theta}^{t},\hat{v}_{i}^{t})$. The algorithm for T-UPBReD is detailed in Algorithm˜3 of the Appendix.
With this mechanism, we can show the following result, the proof of which is deferred to Section˜A.2.
Theorem 2.
Under ˜2, T-UPBReD is EPIC in every round.
In Section˜3.1, we present an example where the NE is not socially optimal, motivating the need for a method to achieve the social optimum as a NE, which we address in the following section.
4 Welfare maximization at Nash equilibrium
In this section, we consider the scenario where the center can make a payment $p_{i}(s_{i},s_{-i})$ to agent $i$ to alter its utility function. Therefore, the effective cost of agent $i$ becomes $z_{i}(s_{i},s_{-i})=c_{i}(s_{i})-p_{i}(s_{i},s_{-i})$. Note that, the payment can be either positive or negative, which determines whether the agent is subsidized or taxed respectively. It is reasonable to expect that the data contributors of the FL process are subsidized for their data contribution while the data consumers are charged payment for obtaining the learned parameter. We assume that the derivatives of the valuation $v_{i}$ and cost function $c_{i}$ are bounded. We choose the following payment function:
$p_{i}(s)=\beta(s_{i}-\frac{1}{n-1}\sum_{j\neq i}s_{j}),$ (5)
where $\beta$ is a parameter of choice. Note that this payment mechanism is budget balanced by design, since $\sum_{i\in N}p_{i}(s_{i},s_{-i})=0,\forall s_{i}\in S_{i},s_{-i}\in S_{-i}$. With payment, the utility of agent $i$ becomes $u_{i}(w,s)=v_{i}(w,s)-c_{i}(s_{i})+p_{i}(s).$ In this section, we are also interested in the quality of the NE in terms of social welfare and want to reach the desired NE where $s_{i}^{*}=s_{i}^{\max},\forall i\in N$ and $w^{*}=w^{\texttt{{OPT}}}$.
4.1 Convergence to the NE $(w^{\textnormal{OPT}},s^{\max})$
In this setup, we first prove that the utility of every agent $i$ can be made strictly increasing in their own contribution $s_{i}$ in the following manner.
Assumption 3.
The derivative of $v_{i}(w,s)$ with respect to $s_{i}$ is bounded away from -$\infty$, i.e., $\frac{\partial}{\partial s_{i}}v_{i}(w,s)\geqslant-\tau$ for some $0\leqslant\tau<\infty$, for all $i\in N,w$, and $s$. Moreover, the cost functions $c_{i}(.)$ has bounded derivatives, i.e., $c^{\prime}_{i}(s_{i})\leqslant\zeta$, for all $s_{i}$ and $i$.
Note that the above assumptions are pretty mild. Previous works assume stronger assumptions (like convexity) on $v_{i}$ and $c_{i}$ (see Murhekar et al. (2023)). We show that the above bounds on the derivatives of $v_{i}$ and $c_{i}$ are sufficient to claim our result.
Lemma 1 (Increasing utility).
Suppose Assumption 3 holds and we use the payment function given by Equation˜5. Then, the utility of every agent is increasing in its own data contribution, i.e., $\frac{\partial}{\partial s_{i}}u_{i}(w,s_{i},s_{-i})>0,\forall s_{i}\in(0,s_{i}^{\max}),\forall i\in N,\text{ and }\forall w$ provided $\beta>\zeta+\tau$.
Remark (Knowledge of $\beta$). The knowledge of $\beta$ is required for theoretical tractability only. In the experiments, we do not assume any knowledge of $\beta$, rather we tune this hyperparameter. We observe that uniformly across all datasets, if $\beta$ is chosen larger than a threshold, we obtain the best performance of our proposed algorithms, which also validates the above theoretical requirement.
This lemma implies that even in Algorithm˜1 if we apply the above payment, $s$ will converge to the maximum value $s^{\max}$. However, unlike Algorithm˜1, in this section we provide an algorithm which gives the step sizes of the gradient ascents in a more concrete manner and is, therefore, a superior one. In order to keep the convergence rate same as Theorem˜1, we assume that the negative [^2] We negate the welfare so that we can consider functions as convex to apply the results of convex analysis easily. social welfare function at $s^{\max}$, given by $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$, is $M$-smooth and $\nu$-strictly convex in $w$. These properties are formally defined below.
Definition 4 ($M$-smoothness).
A function $f:\mathbb{R}^{m}\to\mathbb{R}$ is $M$-smooth if for all $x,x^{\prime}\in\mathbb{R}^{m}$, we have $f(x^{\prime})\leqslant f(x)+\langle\nabla_{x}f(x),x^{\prime}-x\rangle+\frac{M}{2}\|x-x^{\prime}\|_{2}^{2}$.
Definition 5 ($\nu$-strictly convex).
A function $f:\mathbb{R}^{m}\to\mathbb{R}$ is $\nu$-strictly convex if for all $x,x^{\prime}\in\mathbb{R}^{m}$, we have $f(x^{\prime})\geqslant f(x)+\langle\nabla_{x}f(x),x^{\prime}-x\rangle+\frac{\nu}{2}\|x-x^{\prime}\|_{2}^{2}$.
1
2
Input: Step size $\gamma,\eta$, initialization $w^{0}$, $s_{i}^{0}$ for $i\in N$, number of iterations $T$
Output: $w^{T}$
3 Center broadcasts $w^{0},s^{0}$, set $\tau=0$
4while $s\neq s^{\max}$ do
5 Center broadcasts $s^{\tau}$
6 for agent $i\in N$ in parallel do
7    $s_{i}^{\tau+1}=s_{i}^{\tau}+\gamma[g(w^{0},s^{\tau},\mu^{\tau})]_{i}$, send $s_{i}^{\tau+1}$ to center
8
9 $\tau=\tau+1$
10for $t=0$ to $T$ do
11 Center broadcasts $w^{t}$
12 for agent $i\in N$ in parallel do
13      Compute $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
14      Send $d_{i}^{t+1}$ to the center
15
16 Center updates: $w^{t+1}=w^{t}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t+1}=w^{t}+\eta\tilde{g}(w^{t},s^{\max})$
17return $w^{T}$
Algorithm 2 2P-UPBReD
We propose a two-phase algorithm given by Algorithm˜2. The algorithm, in the first phase, incentivizes the agents to contribute $s^{\max}$, and in the second phase, converges to $w^{\texttt{{OPT}}}$. Note that in the first phase of Algorithm˜2, every agent runs a gradient ascent step. Thanks to the increasing utility property (Lemma 1), we have an increasing sequence of $s_{i}^{t}$ for all $i\in N$. Moreover, from the definition of $g(w^{0},s^{t},\mu^{t})]_{i}$ it is ensured that $s_{i}^{\max}$ is the fixed point of this gradient ascent update. Hence after a finite number of iterations, every agent reaches $s_{i}^{\max}$. Note that in this phase the model parameter $w^{0}$ remains unchanged.
In the second phase of the algorithm, we update the model parameter $w^{t}$. Note that since all the agents contribute the maximum amount of data they own, this phase is simply pure FL (without incentive design). As such, each agent now computes the local gradient $d_{i}^{t+1}$ and sends it to the center in each round.
We now provide the convergence guarantees of Algorithm˜2. Before that, let us discuss the necessary assumptions.
Assumption 4.
The negative social welfare function at $s^{\max}$, $f(w,s^{\max})$ , is $M$-smooth and $\nu$-strictly convex in $w$, with $M>\nu$.
The smoothness and strong convexity assumptions have featured in several previous papers on FL Karimireddy et al. (2020b); Yin et al. (2018). The main result of this section is as follows.
Theorem 3 (2P-UPBReD convergence to the optimal NE).
Suppose Assumptions  3 and 4 hold and we consider the utility function with payment scheme given in Equation˜5 with $\beta>\zeta+\tau$. Then, for every $w^{0}\in\mathbb{R}^{m}$ and $s^{0}\in\prod_{i\in N}S_{i}$, 2P-UPBReD (Algorithm˜2) converges to the Nash equilibrium $(w^{\texttt{{OPT}}},s^{\max})$ and is budget balanced. In particular, for any given $\epsilon>0$, we have $\frac{1}{n}\sum_{i\in N}v_{i}(w^{\texttt{{OPT}}},s^{\max})-\frac{1}{n}\sum_{i\in N}v_{i}(w^{T},s^{\max})<\epsilon$, and $s^{T}=s^{\max}$ in $T=\kappa+T_{0}$ iterations provided we choose $\gamma=c$ (an universal constant) and $\eta=1/M$ with
$\kappa\geqslant\max_{i}\left\{\frac{(s^{\max}_{i}-s^{0}_{i})}{c\,\Delta}\right\},$ (6)
and $T_{0}>\left(\ln{\frac{f(w^{0},s^{\max})-f(w^{\texttt{{OPT}}},s^{\max})}{\epsilon}}\right)\Big/\left(\ln\frac{1}{{1-\frac{\nu}{M}}}\right)$ where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s)$ and $\Delta=\beta-\tau-\zeta$.
Remark (Strong convexity and smoothness). The convergence rate is $\mathcal{O}(\ln{\nicefrac{{1}}{{\epsilon}}})$ which is the same as that of Algorithm˜1. However, Algorithm˜2 is cleaner in terms of the phases of convergence, the choices of the step sizes $\gamma,\eta$. The structural assumptions like strong convexity and smoothness may be relaxed at the expense of weaker rates. For example, we observe convergence rates of $\mathcal{O}(\nicefrac{{1}}{{\sqrt{\epsilon}}})$ for non-convex smooth functions and $\mathcal{O}(\nicefrac{{1}}{{\epsilon}})$ for convex smooth functions.
Remark ($\gamma$ and rationality). Phase 1 of Algorithm˜2 achieves the optimal contribution $s^{\max}$ without updating the learning parameter $w$. The speed of convergence of this phase depends on the parameter $\gamma=c$, which can be interpreted as the minimum rationality level of the society. For instance, if all agents are sufficiently rational, i.e., $c$ is large, they may converge to $s^{\max}$ in a single iteration. However, the algorithm converges even if agents are boundedly rational. In Phase 2 of the algorithm, the agents stop updating their contributions and only update the gradients $d_{i}$ and the center accumulates them to update $w^{t}$.
Thanks to the smoothness and strong convexity properties of $f(\cdot,s^{\max})$, we can show that running Algorithm˜2 guarantees that the model iterate $w^{T}$ converges to $w^{\texttt{{OPT}}}$ at an exponential speed as well.
Corollary 1 (Iterate Convergence).
With the same setup as above and $\gamma=c$ (universal constant), $\eta=\frac{2}{M+\nu}$, we obtain $\|w^{T}-w^{\texttt{{OPT}}}\|_{2}<\epsilon,\text{ and }s^{T}=s^{\max}$, where $T=\kappa+\tilde{T}_{0}$, with the same $\kappa$ as in Equation˜6 and $\tilde{T}_{0}>\left(\ln{\frac{\|w^{0}-w^{\texttt{{OPT}}}\|_{2}}{\epsilon}}\right)/\left(\ln\frac{{1+\frac{\nu}{M}}}{{1-\frac{\nu}{M}}}\right)$.
Remark (Truthfulness). An agent’s true type is $(s_{i}^{max},d_{i}^{t})$ in every iteration of phase $2$. We can use the payment scheme $p_{i}^{*}$ as defined in Equation˜4, with the necessary assumptions, in addition to the $z_{i}(s_{i},s_{-i})$ defined in this section, making the mechanism truthful with the trade-off that it will no longer be budget-balanced. Green and Laffont (1979) show that it is impossible to achieve BB, truthfulness and optimality together.
Remark (Robustness to adversarial behavior). In 2P-UPBReD, we have assumed that the agents report their gradients truthfully in phase $2$. If the agents misreport by sending arbitrary gradients not computed on any data, we have discussed one option of enforcing additional payment in the previous remark that compromises budget balance. If we do not wish to enforce additional payments, we can use some established methods in literature that are robust to adversarial agents (Yin et al., 2018; Karimireddy et al., 2020a; Ghosh et al., 2021). We describe one such method called trimmed mean  (Yin et al., 2018), in Section˜B.4 and implement 2P-UPBReD with trimmed mean in the following experiments
In Section˜B.4, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In this section, we demonstrate this robustness on the FEMNIST and CIFAR-10 dataset. We vary the adversarial fraction $\alpha\in\{0.025,0.05,0.075,0.1,0.125,0.15,0.175,0.2\}$, selecting suitable trimming parameters. Figure˜2 shows the ratio of social welfare with and without adversaries for UPBReD and 2P-UPBReD. While both suffer welfare loss under adversaries, applying the trimmed mean algorithm (Algorithm˜4) keeps the welfare ratio close to $1$ for trimmed mean 2P-UPBReD.
(a) CIFAR-10
(b) FEMNIST
Figure 2: Ratio of social welfare with adversaries and without adversaries.
5 Experiments
Our results in Sections˜3 and 4 guarantee convergence of Algorithms˜1 and 2 to a pure Nash equilibrium under the given assumptions. However, it is important to ask how these FL algorithms behave in practice with real datasets under more general setups. We capture results in these realistic settings in our experiments. We train models on the CIFAR-10  Krizhevsky (2009), FEMNIST, and Twitter datasets Caldas et al. (2019) and compare the social welfare of UPBReD, 2P-UPBReD, and FedAvg. We also consider another algorithm FedAvgStrategic as a baseline. In this algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg (details in Section˜C.1). The mechanisms/algorithms proposed in Table˜1, which are closest to this paper, do not use valuation functions that consider simultaneous learning and strategy updates and are therefore not meaningful to compare against. The cost function used is linear $c_{i}(s_{i})=c_{i}\cdot s_{i}$, where $c_{i}$ is sampled from $U[0,1]$ for each agent. We choose our valuation function to be $v_{i}(w^{k},s^{k})=r_{i}-\mathcal{L}_{i}(w^{k},s^{k})$ where the $\mathcal{L}_{i}(w^{k},s^{k})$ is the loss evaluated on agents $i$’s test dataset at the $k$-th round of training. The social welfare is given by $\sum_{i\in N}v_{i}(w^{k},s^{k})$. We perform strategy updates using the numerical derivative as follows: $s_{i}^{k+1}=s_{i}^{k}-\frac{L_{i}(w_{i}^{k})-L_{i}(w^{k-1})}{s_{i}^{k}-s_{i}^{k-1}}-c_{i}+\beta$.
The model parameters obtained by agent $i$ after performing one training step locally on the model initialized with $w^{k-1}$ is denoted by $w_{i}^{k}$. We use the cross-entropy loss function for $L_{i}$} on agent i’s dataset. All the experiments are performed on NVIDIA RTX A5000 and NVIDIA RTX A6000 with a 24GB and 48GB core respectively. The experiments required to obtain the results we report take several days to run. For comparing social welfare against $\beta$ and number of agents, we randomize over 10 runs, sampling an integer $s^{0}_{i}$ from $U[\frac{s^{\max}_{i}}{3},\frac{2s^{\max}_{i}}{3}]$ and cost from $U[0,1]$. We plot the average value of time taken across $10$ runs.
Dataset specifics. For CIFAR-10, the data is distributed equally among 100 agents by sampling the data uniformly at random. We train a CNN with 1,250,858 parameters for 100 rounds. Hyperparameters: $\gamma=0.5,\ \beta=2,\ \eta=0.001$, optimizer=Adam.
For FEMNIST, we use $20\%$ of the dataset that generates users with data that is not identically distributed as per the codebase provided by Caldas et al. (2019). We train a CNN with $6,603,710$ parameters for $100$ rounds. Hyperparameters: $\gamma=0.1,\ \beta=2,\ \eta=0.001$, optimizer=Adam.
For Twitter, we use $15\%$ of the dataset that generates users with data that is identically distributed. We further group $1000$ agents together to create agents with larger data shares. We use an LSTM network to train for $50$ rounds. Hyperparameters: $\gamma=0.5,\ \beta=2,\ \eta=0.00001$, optimizer=Adam.
All results are summarized in Figure˜3. We provide more details of the model architecture for all models in Section˜C.2.
(a) CIFAR-10
(b) FEMNIST
(c) Twitter
Figure 3: The three rows in this figure correspond to the experiments with the datasets CIFAR-10, FEMNIST, and Twitter, respectively. The y-axis in every plot is the social welfare. The x-axis corresponds to the number of agents in the first plot, the choice of $\beta$ in the second, and the time taken (in seconds) in the third plot, respectively, in every row.
We observe that 2P-UPBReD and FedAvg have the same social welfare performance w.r.t. the number of agents (first column), and running times are nearly the same for all the datasets (third column). 2P-UPBReD outperforms UPBReD in nearly all the runs with respect to social welfare. We also observe that 2P-UPBReD converges to the maximum contribution of the agents for values of $\beta$ that are lower than the theoretically predicted ones. We also note that the data contributors get paid and data consumers pay in 2P-UPBReD and their payments are monotone non-decreasing with data contribution (Figure˜4). The utilities vary because the dataset on which the model is tested to obtain the accuracy varies for the agents.
Figure 4: Utilities and payments received by agents for 2P-UPBReD based on their data contributions for FEMNIST (averaged over 10 runs).
We leverage the partitioners provided by the Flower framework to generate multiple data splits of the CIFAR-10 dataset. We evaluate the performance of the algorithms under three data-partitioning regimes: an IID split, a Dirichlet split, and a Pathological split. The resulting distribution for the Dirichlet and Pathological split is visualised in Figure˜6. For the Dirichlet split, the concentration parameter is set to $\alpha=0.1$ with self-balancing enabled. For the Pathological split, the class assignment mode is set to random, and each partition contains $3$ classes. Hyperparameter details: $\beta=2,\gamma=1,\eta=0.0001$,optimizer=Adam. The results are summarized in Figure˜5, averaged across $5$ runs.
In Section˜B.4, we describe the trimmed mean algorithm for 2P-UPBReD designed to be robust in the presence of adversarial agents. In LABEL:app:experiments_adversarial we show that 2P-UPBReD with trimmed mean is robust to adversarial agents.
(a) IID Split
(b) Dirichlet Split
(c) Pathological Split
Figure 5: Social welfare vs. number of clients for different CIFAR-10 dataset splits using the Flower framework.
(a) Pathological partitioner (33 classes per client).
(b) Dirichlet partitioner ($\alpha=0.1$).
Figure 6: Label distributions for different data partitioning strategies using Flower with $10$ agents. The x-axis represent the client id and the y-axis represents the number of data points of the given label.
6 Conclusions and future work
In this paper, we proposed UPBReD, which ensures convergence to a Nash equilibrium with model parameters $w$, and its modification T-UPBReD, which enforces truthful reporting. Since $w$ may deviate from the optimal $w^{\texttt{{OPT}}}$, we introduced a two-phase mechanism, 2P-UPBReD, that guarantees full agent participation and convergence to $w^{\texttt{{OPT}}}$, at the cost of monetary transfers that can be internal among agents. Future directions include incorporating agent budget constraints to preserve approximate guarantees and extending the framework to heterogeneous data settings, enabling clustered learning of optimal parameters.
Appendix A Appendix to Section˜3
A.1 Proof of Theorem˜1
Consider the first-order Taylor expansions of $g$ and $\tilde{g}$ given as follows.
$g(w^{t+1},s^{t+1},\mu^{t+1})$ $=g(w^{t},s^{t},\mu^{t})+G(w^{t},s^{\prime},\mu^{t})\cdot(s^{t+1}-s^{t})$
$\quad+{H(w^{\prime},s^{t},\mu^{t})(w^{t+1}-w^{t})},$
$\tilde{g}(w^{t+1},s^{t+1})$ $=\tilde{g}(w^{t},s^{t})+\tilde{G}(w^{\prime},s^{t})\cdot(w^{t+1}-w^{t})$
$\quad+\tilde{H}(w^{t},s^{\prime})(s^{t+1}-s^{t}),$
where $s^{\prime}=\theta_{s}s^{t}+(1-\theta_{s})s^{t+1},w^{\prime}=\theta_{w}w^{t}+(1-\theta_{w})w^{t+1}$, for some $\theta_{s},\theta_{w}\in[0,1]$.
Using the update rules of Algorithm˜1 given by Equation˜3, we get
$g(w^{t+1},s^{t+1},\mu^{t+1})$ $=g(w^{t},s^{t},\mu^{t})+G(w^{t},s^{\prime},\mu^{t})\cdot\gamma g(w^{t},s^{t},\mu^{t})$
$\quad+{H(w^{\prime},s^{t},\mu^{t})\eta\tilde{g}(w^{t},s^{t})}$
$\tilde{g}(w^{t+1},s^{t+1})$ $=\tilde{g}(w^{t},s^{t})+\tilde{G}(w^{\prime},s^{t})\cdot\eta\tilde{g}(w^{t},s^{t})$
$\quad+\tilde{H}(w^{t},s^{\prime})\gamma g(w^{t},s^{t},\mu^{t})$
Using triangle inequality on each of these identities, we get
$\begin{aligned} \textstyle\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}&\textstyle\leqslant\|(I_{n\times n}+\gamma G(w^{t},s^{\prime},\mu^{t}))g(w^{t},s^{t},\mu^{t})\|_{2}\\ \textstyle&\textstyle\quad+\eta\|{H(w^{\prime},s^{t},\mu^{t})\tilde{g}(w^{t},s^{t})}\|_{2}\\ \textstyle\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}&\textstyle\leqslant\|(I_{m\times m}+\eta\tilde{G}(w^{\prime},s^{t}))\tilde{g}(w^{t},s^{t})\|_{2}\\ \textstyle&\textstyle\quad+\gamma\|\tilde{H}(w^{t},s^{\prime})g(w^{t},s^{t},\mu^{t})\|_{2}\end{aligned}$ (7)
From condition 1 of ˜1, we get that $v^{\top}(G+\lambda I_{n\times n})v\leqslant 0,\forall v\in\mathbb{R}^{n}$ and $v^{\prime\top}(\tilde{G}+\lambda I_{m\times m})v^{\prime}\leqslant 0,\forall v^{\prime}\in\mathbb{R}^{m}$. In particular, for $v=g(w^{t},s^{t},\mu^{t})$ and $v^{\prime}=\tilde{g}(w^{t},s^{t})$, we get
$\begin{split}\textstyle g(w^{t},s^{t},\mu^{t})^{\top}G(w^{t},s^{\prime},\mu^{t})g(w^{t},s^{t},\mu^{t})&\textstyle\leqslant-\lambda\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\\ \textstyle\tilde{g}(w^{t},s^{t})^{\top}\tilde{G}(w^{\prime},s^{t})\tilde{g}(w^{t},s^{t})&\textstyle\leqslant-\tilde{\lambda}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}\end{split}$ (8)
Consider the square of the first term of the RHS of the inequality of $g$ in Equation˜7
[…]
where the equality comes by expanding the squared norm and the inequality comes from the facts that (i) $\|G(w^{t},s^{t})g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\leqslant\|G(w^{t},s^{t})\|_{F}^{2}\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}$, where $\|A\|_{F}:=\sqrt{\sum_{i}\sum_{j}|A_{ij}|^{2}}$ is the Frobenius norm of a matrix $A$ and (ii) using Equation˜8. By condition 2 of ˜1, $\|G(w^{t},s^{t})\|_{F}^{2}\leqslant n^{2}L^{2}$. Hence, we get
$\textstyle\|(I_{n\times n}+\gamma.G(w^{t},s^{t}))g(w^{t},s^{t},\mu^{t})\|$
$\textstyle\leqslant\sqrt{1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda}\cdot\|g(w^{t},s^{t},\mu^{t})\|_{2}$
given the term inside the square root is positive.
Consider the second term of the RHS of the inequality of $g$ in Equation˜7, where
$\textstyle\|H(w^{\prime},s^{t},\mu^{t})\tilde{g}(w^{t},s^{t})\|_{2}^{2}$ $\textstyle\leqslant\|H(w^{\prime},s^{t},\mu^{t})\|_{op}^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
$\textstyle\leqslant P^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
by condition 2 of ˜1.
Defining $\alpha:=1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda,\beta:=P^{2}\eta^{2},\tilde{\alpha}:=1+\eta^{2}.m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda},\tilde{\beta}:=\tilde{P}^{2}\gamma^{2}$, and carrying out a similar analysis for $\tilde{g}$ in Equation˜7, we get
$\begin{split}\textstyle\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}&\textstyle\leqslant\sqrt{\alpha}\|g(w^{t},s^{t},\mu^{t})\|_{2}+\sqrt{\beta}\|\tilde{g}(w^{t},s^{t})\|_{2}\\ \textstyle\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}&\textstyle\leqslant\sqrt{\tilde{\alpha}}\|\tilde{g}(w^{t},s^{t})\|_{2}+\sqrt{\tilde{\beta}}\|g(w^{t},s^{t},\mu^{t})\|_{2}\end{split}$ (10)
Adding the inequalities of Equation˜10
[…]
To ensure that the above inequality is a contraction, we need to ensure $\sqrt{\alpha}+\sqrt{\tilde{\beta}},\sqrt{\tilde{\alpha}}+\sqrt{\beta}\in(0,1)$. These imply (i) $0<\alpha<1,0<\tilde{\alpha}<1$, (ii) $\sqrt{\tilde{\beta}}<1,\sqrt{\beta}<1$, and (iii) $\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1,\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$. We can solve for $\gamma$ and $\eta$ from these inequalities and obtain the sufficient conditions when $\lambda>\tilde{P}$ and $\tilde{\lambda}>P$:
$\gamma$ $<\frac{1}{\tilde{P}},\gamma<\frac{2\lambda}{n^{2}L^{2}},\text{ and }\gamma<\frac{\lambda-\tilde{P}}{n^{2}L^{2}-\tilde{P}^{2}},$
$\eta$ $<\frac{1}{P},\eta<\frac{2\tilde{\lambda}}{m^{2}\tilde{L}^{2}},\text{ and }\eta<\frac{\tilde{\lambda}-P}{m^{2}\tilde{L}^{2}-P^{2}}.$
Notice that from conditions 1 and 2 of ˜1, we have that $\lambda<L$ and $\tilde{\lambda}<\tilde{L}$. Thus $n^{2}L^{2}-\tilde{P}^{2}>0$ and $m^{2}\tilde{L}^{2}-P^{2}>0$. For the $\gamma,\eta$ chosen as above, we find $W_{1}=\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1$ and $W_{2}=\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$, and hence $W=\max\{W_{1},W_{2}\}<1$. Therefore,
$\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}+\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}$
$\leqslant W\left(\|{g}(w^{t},s^{t},\mu^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}\right).$
Recursively iterating over this inequality, we get
$\textstyle\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}$
$\textstyle\leqslant W^{T}\left(\|{g}(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}\right).$
Defining $E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}$ and
$T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\textstyle E}{\textstyle\epsilon}}\right)}}{{\left(\ln{\frac{\textstyle 1}{\textstyle W}}\right)}},$
we get that for all $T\geqslant T_{0}(w^{0},s^{0})$,
$\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}<\epsilon$
$\implies$ $\|g(w^{T},s^{T},\mu^{T})\|_{2}<\epsilon,\text{ and }\|\tilde{g}(w^{T},s^{T})\|_{2}<\epsilon.$
This completes the proof.
A.2 Omitted examples and proofs of Section˜3.2
We perform some experiments on the CIFAR data set to show that agents have an incentive to misreport. Consider $2$ agents each having a set of $480$ data points. $u_{i}=56-L(w,s)-c_{i}(s_{i})$, where $L(w,s)$ is the cross entropy loss computed on the dataset. The costs for the agents are $c_{1}(s_{1})=0.2s_{i},c_{2}(s_{2})=0.1s_{2}$. When agents truthfully report $s_{i},d_{i}$ in UPBReD they observe a social welfare of $15.5$, with $s_{1}*=270,s_{2}*=165$ this yields a utility $u_{1}=510$. When agent misreports $s_{1}^{t}=240$ and reports $d_{1}^{t}=0_{m}$ in every training round yielding a utility of $u_{i}=64$. Clearly agent $1$ is better off by misreporting their type.
The algorithm to implement T-UPBReD is described below.
1
Input: Step size $\gamma,\eta$, initialization $w^{0}$, $s_{i}^{0}$ for $i\in N$, number of iterations $T$
Output: $w^{T}$
2
3for $t=0$ to T1T-1 do
4 Center broadcasts $w^{t},s^{t}$
5
6 foreach agent $i\in N$ in parallel do
7    $s^{t+1}_{i}=s^{t}_{i}+\gamma[g(w^{t},s^{t},\mu^{t})]_{i}$
8      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{t},s_{-i}^{t})$
9      Send $\tilde{s}_{i}^{t+1},\tilde{d}_{i}^{t+1}$ to the center
10
11
12 Center updates: $w^{t+1}=w^{t}+\frac{\eta}{n}\sum_{i\in N}\tilde{d}_{i}^{t+1}$
13
14 Center broadcasts $w^{t+1}$
15 foreach agent $i\in N$ in parallel do
16      Observes $v_{i}(w^{t+1},\tilde{s}^{t+1})$
17     Send $\hat{v}_{i}^{t}=\hat{a}_{i}$
18
19 Center computes payments using Equation˜4 at $w^{t}$
20
21return $w^{T}$
22
Algorithm 3 Truthful Updated Parameter Best Response Dynamics (T-UPBReD)
Proof of Theorem˜2
Consider the utility of agent $i$ at time $t$, whose true type and valuation are $\theta_{i}^{t}=(s_{i}^{t},d_{i}^{t}),v_{i}^{t}$ respectively, misreporting their type and valuation, $\hat{\theta}^{t}_{-i}=(\hat{s}_{i}^{t},\hat{d}_{i}^{t})$, $\hat{v}_{i}^{t}$ while other agents are truthful $\theta_{-i^{t}}=(s_{-i}^{t},d_{-i}^{t})$.
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$=v_{i}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})+p_{i}(\hat{\theta}^{t}_{i},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t})$
Expanding the terms as defined in Equation˜4
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$=v_{i}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})+\sum_{j\neq i}v_{j}(a_{w}^{t}(\hat{\theta}_{i}^{t},\theta_{-i}^{t}),s^{t})$
$-\Gamma-(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{\theta}^{t}))^{2}-h_{i}(\theta_{-i}^{t})$
Since $(\hat{v}^{t}_{i}-v_{i}(a_{w}^{t}(\hat{\theta}^{t}),\hat{\theta}^{t}))^{2}$ is always non-positive, we get
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leq\sum_{j\in N}v_{j}(a_{w}^{t}(\hat{\theta}^{t}_{i},\theta_{-i}^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-\Gamma-h_{i}(\theta_{-i}^{t})$
As social welfare is maximized by $a_{*}(\theta^{t})$
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leqslant\sum_{j\in N}v_{j}(a_{*}(\theta^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-\Gamma-h_{i}(\theta_{-i}^{t})$
Using ˜2
$u_{i}(\hat{\theta}_{i}^{t},\theta_{-i}^{t},\hat{v}_{i}^{t},v_{-i}^{t}|\theta^{t},v^{t})$
$\leqslant\sum_{j\in N}v_{j}(a_{w}^{t}(\theta^{t}),s^{t})+\Gamma-\Gamma-z_{i}(s_{i}^{t},s_{-i}^{t})-h_{i}(\theta_{-i}^{t})$
$=\sum_{j\in N}v_{j}(a_{w}^{t}(\theta^{t}),s^{t})-z_{i}(s_{i}^{t},s_{-i}^{t})-h_{i}(\theta_{-i}^{t})$
$=u_{i}(\theta^{t},v^{t}|\theta^{t},v^{t})$
Appendix B Appendix to Section˜4
B.1 Proof of Lemma˜1
Proof.
Using Equation˜5, we get the derivative of the utility function for $s_{i}\in(0,s_{i}^{\max})$ (note that this holds only for the interior of $S_{i}$, the derivative at the boundaries are zero by definition of function $g$ in Section˜3) to be $g(w,s,\mu)]_{i}=\frac{\partial}{\partial s_{i}}u_{i}(w,s)=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta$, where $c_{i}^{\prime}(s_{i})=\frac{d}{ds_{i}}c_{i}(s_{i})$. Using Assumption 3 and choosing in Equation˜5 the parameter $\beta>\zeta+\tau$, we have
$[g(w,s,\mu)]_{i}$ $=\frac{\partial}{\partial s_{i}}v_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta\geqslant-\tau-\zeta+\beta>0.$
The last inequality holds since $\beta>\zeta+\tau$, $\forall i\in N$. ∎
B.2 Proof of Theorem˜3
Proof.
Phase 1: Using 1
$\begin{split}\textstyle[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant-\tau-\zeta+\beta>0\end{split}$ (12)
For the choice of $\beta>\zeta+\tau$ , we get $\Delta:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm˜2 in phase 1, we get
$s^{t+1}_{i}$ $=s^{t}_{i}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant s^{t}_{i}+\gamma\Delta,$
and applying the inequality repeatedly for $t$ iterations yields
$s^{t}_{i}\geqslant s^{0}_{i}+\gamma(t\Delta).$
Let $l_{i}=s^{\max}_{i}-s^{0}_{i}$. We have,
$s^{t}_{i}$ $\geqslant s^{\max}_{i}-l_{i}+\gamma(t\Delta)$
Substituting $t=\kappa$ where $\kappa\geqslant\frac{\textstyle l_{i}}{\textstyle\Delta\gamma}$, we obtain
$\begin{split}\textstyle s^{\kappa}_{i}&\textstyle\geqslant s^{\max}_{i}-l_{i}+\gamma\Delta(\frac{\textstyle l_{i}}{\textstyle\Delta\gamma})\\ \textstyle&\textstyle\geqslant s^{\max}_{i}-l_{i}+l_{i}\\ \textstyle&\textstyle\geqslant s^{\max}_{i}\\ \textstyle s_{i}^{\kappa}&\textstyle\geqslant s_{i}^{\max}\end{split}$ (13)
Since, $s_{i}^{t}\leqslant s^{\max}_{i}$ for all $t$ and $i\in N$, we conclude that $s^{\kappa}_{i}=s^{\max}_{i}$ for all $i\in N$. Hence, $s^{\max}$ is a fixed point of the update $s^{t+1}=s^{t}+\gamma g(w^{0},s^{t},\mu^{t})$ at the end of phase 1.
Phase 2: For $t>\kappa$, we can focus on the second phase of the algorithm. From the fixed point property of $s^{\max}$, we can now focus entirely on the function $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$. From the algorithmic description, the second phase is just a simple gradient descent, run by the center, on $f(w,s^{\max})$. This is easy to see since at every iteration the center gets $d_{i}^{t+1}$ from all the agents $i\in N$, aggregates them and construct $\tilde{g}(w^{t},s^{\max})$, which is the gradient of $f(w,s^{\max})$ computed at $w^{t}$. We denote $f(w,s^{\max})$ with $f(w)$ in this part for notational cleanliness.
Note that the gradient descent is run with initialization $w^{\kappa}$ (which is same as $w^{0}$ as given by first phase of the algorithm). We exploit the strong convexity and smoothness of $f(w,s^{\max})$. Running the second phase for $T_{0}$ iterations, using Wright and Recht (2022), we obtain
$f(w^{T_{0}})-f(w^{\texttt{{OPT}}})\leqslant(1-\frac{\nu}{M})^{T_{0}}[f(w^{0})-f(w^{\texttt{{OPT}}})].$
Hence, for $f(w^{T_{0}})-f(w^{\texttt{{OPT}}})<\epsilon$, we require
$T_{0}>\left(\ln{\frac{f(w^{0})-f(w^{\texttt{{OPT}}})}{\epsilon}}\right)\Big/\left(\ln\frac{1}{\Big({1-\frac{\nu}{M}\Big)}}\right),$
which proves the theorem. ∎
B.3 Proof of Corollary˜1
Proof.
The proof of this follows in the same lines as Theorem 3. Using the choice of $\beta$, we first ensure that with $\kappa$ steps, we obtain $s_{i}^{\kappa}=s^{\max}_{i}$ for all $i\in N$. Now, the framework is same as minimization of a strongly convex and smooth function $f(w)$ with initialization $w^{\kappa}=w^{0}$. Using Wright and Recht (2022), with $\eta=\frac{2}{M+\nu}$, we obtain the iterate convergence, namely $\|w^{\tilde{T}_{0}}-w^{\texttt{{OPT}}}\|_{2}\leqslant\left(\frac{1-\frac{\nu}{M}}{1+\frac{\nu}{M}}\right)^{\tilde{T}_{0}}\|w^{0}-w^{\texttt{{OPT}}}\|_{2}$. Taking log both sides implies the result. ∎
B.4 Adversarial agents
Our results for 2P-UPBReD assume that agents are honest and report their gradients truthfully. However, the agents may be adversaries who misreport their gradients. Prior works (Yin et al., 2018) describe some simple ways to deal with this. When the ratio of adversarial agents in the system is low we use 2P-UPBReD along with a trimmed mean algorithm described in (Yin et al., 2018).
Let $\alpha$ be the fraction of adversarial agents in the system. Phase $1$ of the algorithm remains the same with agents first converging to $s^{\max}$. In the second phase which is a classic FedAvg algorithm the agents only report gradients computed on their training data. Instead of aggregating all gradients, for each coordinate of the gradients the values of the gradient corresponding to that coordinate are sorted. The lowest and highest $\alpha$ fraction of these values are discarded and the remaining gradient coordinates are used to update the parameter. We detail the algorithm in Algorithm˜4. In LABEL:app:experiments_adversarial we show that this algorithm is robust to adversaries when the fraction of adversarial agents is small.
1
[…]
2
Output: $w^{T}$
3 Center broadcasts $w^{0},s^{0}$, set $t=0$ while $s\neq s^{\max}$ do
4 Center broadcasts $s^{t}$
5 for agent $i\in N$ in parallel do
6    $s_{i}^{t+1}=s_{i}^{t}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}$
7      Send $s_{i}^{t+1}$ to the center
8
9 $t=t+1$
10for $t=0$ to $T$ do
11 for agent $i\in N$ in parallel do
12      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
13      Send $d_{i}^{t+1}$ to the center
14
15
16 foreach $j\in[m]$ do
17      Construct $W_{j}=\{d_{ij}|i\in N\}$.
18      Sort $W_{j}$ into an ascending array $C$.
19    $C=C[n\alpha,n-n\alpha]$
20    $W_{j}=\{d_{ij}|d_{ij}\in C\}$
21    Center updates: $w^{t+1}_{j}=w^{t}_{j}+\frac{\eta}{n}\sum_{d_{ij}\in W_{j}}\tilde{d}_{ij}^{t+1}$
22
23 Center broadcasts $w^{t+1}=w$
24
return $w^{T}$
Algorithm 4 Trimmed mean for 2P-UPBReD (trim-2P-UPBReD)
Appendix C Appendix to Section˜5
C.1 Algorithm for FedAvgStrategic
In the FedAvgStrategic algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg. FedAvgStrategic is formally given by Algorithm˜5.
1
2
Input: Step size $\gamma,\eta$, initialization $w^{0}$, $s_{i}^{0}$ for $i\in N$, number of iterations $T$
Output: $w^{T}$
Center broadcasts $w^{0},s^{0}$, set $t=0$
/* Begin phase 1 */
3 while $\exists i\in N\text{ such that }[g(w^{0},s^{t},\mu^{t})]_{i}>0$ do
4 Center broadcasts $s^{t}$
5 for agent $i\in N$ in parallel do
6      Perform local training step and compute $g(w^{0},s^{t},\mu^{t})]_{i}$
7    $s_{i}^{t+1}=s_{i}^{t}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}$
8      Send $s_{i}^{t+1}$ to the center
9
 $t=t+1$
 /* End phase 1 */
10
[…]
/* Begin phase 2 */
12 for $t=0$ to TTdo
13 Center broadcasts $w^{t}$
14 for agent $i\in N$ in parallel do
15      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{*},s_{-i}^{*})$
16      Send $d_{i}^{t+1}$ to the center
17
18 Center updates: $w^{t+1}=w^{t}+\frac{\eta}{n}\sum_{i\in N}d_{i}^{t+1}=w^{t}+\eta\tilde{g}(w^{t},s^{*})$
return $w^{T}$
/* End phase 2 */
19
Algorithm 5 FedAvgStrategic
C.2 Model Architecture
We provide details of the model architecture used in our experiments in Tables˜2, 3 and 4.
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 32, 32] 896
Conv2d: 1-2 [1, 32, 30, 30] 9,248
MaxPool2d: 1-3 [1, 32, 15, 15]
Conv2d: 1-4 [1, 64, 15, 15] 18,496
Conv2d: 1-5 [1, 64, 13, 13] 36,928
MaxPool2d: 1-6 [1, 64, 6, 6]
Linear: 1-7 [1, 512] 1,180,160
Linear: 1-8 [1, 10] 5,130
Total Params 1,250,858
Trainable Params 1,250,858
Non-trainable Params 0
Total Mult-Adds (MB) 20.83
Input Size (MB) 0.01
Forward/Backward Pass Size (MB) 0.70
Params Size (MB) 5.00
Estimated Total Size (MB) 5.71
Table 2: Layer-wise Summary of CNNModel for CIFAR
Layer (type:depth-idx) Output Shape Param #
CNNModel [1, 10]
Conv2d: 1-1 [1, 32, 28, 28] 832
MaxPool2d: 1-2 [1, 32, 14, 14]
Conv2d: 1-3 [1, 64, 14, 14] 51,264
MaxPool2d: 1-4 [1, 64, 7, 7]
Linear: 1-5 [1, 2048] 6,424,576
Linear: 1-6 [1, 10] 20,490
Total Params 6,497,162
Trainable Params 6,497,162
Non-trainable Params 0
Total Mult-Adds (MB) 17.15
Table 3: Layer-wise Summary of CNNModel used for FEMNIST
Layer (type:depth-idx) Output Shape Param #
Twitter [1, 2]
Embedding: 1-1 [1, 25, 300] 120,000,300
LSTM: 1-2 [1, 25, 100] 241,600
Linear: 1-3 [1, 128] 12,928
Linear: 1-4 [1, 2] 258
Total Parameters 120,255,086
Trainable Parameters 120,255,086
Non-trainable Params 0
Total Mult-Adds (MB) 126.05
Input Size (MB) 0.00
Forward/Backward Pass Size (MB) 0.08
Params Size (MB) 481.02
Estimated Total Size (MB) 481.10
Table 4: Layer-wise Summary of Twitter Model